# FG-CLIP 2: A Bilingual Fine-grained Vision-Language Alignment Model

### v2

@@ Appendix C Dense Prediction Tasks Experimental Details
Appendix D Visualization of Semantic Alignment Capability for Dense Visual Features
@@
We present the alignment capability of FG-CLIP 2 between dense visual features and text in both Chinese and English contexts. The results are shown in Figure A, where warmer colors indicate higher similarity between image regions and the matched text. Compared to the previous version, FG-CLIP 2 supports denser visual feature output and achieves strong bilingual semantic alignment and fine-grained perception capabilities.
Figure A: Visualization of FG-CLIP 2’s dense feature maps and semantic alignment capability in bilingual scenarios.
Appendix E Examples of Chinese Long Caption Image-Text Pairs
@@
In Table C, we provide examples of long caption image-text pairs from the LIT-CN dataset, covering diverse scene categories such as indoor, outdoor, animals, products, and buildings. These captions not only describe fine-grained subject attributes (e.g., appearance, posture, spatial layout), but also detail the surrounding context, reflecting the dataset’s semantic richness and descriptive complexity.
Table C: Examples from LIT-CN.
@@
Appendix F BoxClass-CN Category Schema and Example Explanation
Table D.1 and D.2 provide the complete list of categories in the BoxClass-CN dataset, separated by commas. We further present some examples from BoxClass-CN in Figure  B.
Table D.1: All categories in BoxClass-CN, grouped and displayed by ID (1-300).
@@ Appendix F BoxClass-CN Category Schema and Example Explanation
Table D.2: All categories in BoxClass-CN, grouped and displayed by ID (301-566).

### v3

@@ Appendix F Hyperparameter Ablation Study
Appendix G Visualization of Semantic Alignment Capability for Dense Visual Features
@@
We present the alignment capability of FG-CLIP 2 between dense visual features and text in both Chinese and English contexts. The results are shown in Figure C, where warmer colors indicate higher similarity between image regions and the matched text. The precise localization of high-similarity regions across both languages demonstrates FG-CLIP 2’s strong bilingual semantic alignment and fine-grained perception capabilities.
Appendix H Visual Analysis of the Textual Intra-modal Contrastive Loss
We conduct an embedding visualization comparing text representations with and without the TIC loss. As shown in the visualization at Table E.1 and Table E.2, when TIC is enabled, semantically similar but distinct phrases become better separated in the embedding space. This suggests that TIC indeed enhances semantic separability among fine-grained textual descriptions.
Appendix I Examples of LIT-CN
@@
In Table F, we provide examples of long caption image-text pairs from the LIT-CN dataset, covering diverse scene categories such as indoor, outdoor, animals, products, and buildings. These captions not only describe fine-grained subject attributes (e.g., appearance, posture, spatial layout), but also detail the surrounding context, reflecting the dataset’s semantic richness and descriptive complexity.
Figure C: Visualization of FG-CLIP 2’s dense feature maps and semantic alignment capability in bilingual scenarios.
Table E.1: The Effectiveness of the Textual Intra-modal Contrastive Loss. (Case1 and Case2)
Case1
Case2
Table E.2: The Effectiveness of the Textual Intra-modal Contrastive Loss. (Case3 and Case4)
Case3
Case4
Appendix J BoxClass-CN Category Schema and Example Explanation
Table G.1 and G.2 provide the complete list of categories in the BoxClass-CN dataset, separated by commas. We further present some examples from BoxClass-CN in Figure D.
Table F: Examples from LIT-CN.
@@
Figure D: Examples from BoxClass-CN.
Table G.1: All categories in BoxClass-CN, grouped and displayed by ID (1-300).
@@
Table G.2: All categories in BoxClass-CN, grouped and displayed by ID (301-566).